OpenStudy (anonymous):

PLEASE HELP???? What is the minimum of 2 x^2+2^{5} x^{-2}?

OpenStudy (anonymous):

Let's start by finding the critical points of the function. \[ f(x) = 2x^2 + 2^5x^{-2} \\ f'(x) = 4x -2^6x^{-3} \\ f'(x) = 0 \quad \to \quad 4x -2^6x^{-3} = 0 \\ 4x^4 - 2^6 = 0 \\ 4x ^4 = 2^6 \\ x^4 = \frac{2^6}{4} = \frac{2^6}{2^2} = 2^4 \\ x = \pm\; 2 \\ \] Now for each let's check the second derivative to see if they are minimum or maximum. \[ f''(x) = \bigg[f'(x) \bigg]' = \bigg[4x -2^6x^{-3}\bigg]' = 4 + 3 \cdot 2^{6}x^{-4}\\ \] \(f''(x) > 0 \) Means the slope of \(f(x)\) is increasing at \(x\), and on a critical point it means it's a minimum point. \[ f''(2) = 4 + 3\cdot 2^6 \cdot 2^{-4} = 4 + 3 \cdot2^2 = 16\\ f''(-2) = 4 + 3\cdot 2^6 \cdot (-2)^{-4} = 4 + 3 \cdot2^2 = 16 \] Means... both \(x=2\) and \(x=-2\) are minimum points of \(f(x)\) at value of 16. For reference: http://www.wolframalpha.com/input/?i=y+%3D+2x%C2%B2+%2B+32x%E2%81%BB%C2%B2 Agrees about that as well